I am taking a pharmacology course and am having trouble remembering how to solve proportion questions. ex(2:x::30:600) ? 20: x :: 30= 10?
Answer by venugopalramana(3286)

ex(2:x::30:600) ? 20: x :: 30: 10?
IT IS REALLY SIMPLE..THERE IS ONLY 1 RULE ..THAT IS...
PRODUCT OF EXTREMES = PRODUCT OF MEANS
2*600 = X*30.......OR.................20*10=X*30
1200=30X...............................200=30X
X=1200/30...............................X=200/30
X=40....................................X=20/3